OpenStudy (anonymous):

Math Analysis: Find the indicated roots and express in both polar and rectangular form.Exact answers please! 15. Fourth roots of 8sqr. rt.3 + 8i

OpenStudy (anonymous):

\[8\sqrt{3} + 8i\]

OpenStudy (anonymous):

\[ \tan (\theta) = \frac {8}{8\sqrt 3}=\frac {1}{\sqrt 3}\\ \theta = \frac\pi 6\\ r=\sqrt { 8^2 + (8 \sqrt 3)^2 }=\sqrt{256}=16 \] What are the polar coordinates?

OpenStudy (anonymous):

how did yo get pi/6?

OpenStudy (anonymous):

oh alrite i get it :)

OpenStudy (anonymous):

what do you mean the polar coordinates?

OpenStudy (anonymous):

it would be 16cis pi/6 rite?

OpenStudy (anonymous):

or 16cis30

OpenStudy (anonymous):

\[ 8\sqrt 3 + 8 i = 16 e^{ i \frac \pi 6} \] How can you write the 4th root of that in polar?

OpenStudy (anonymous):

\[ (16 e^{ i \frac \pi 6})^\frac 1{ 4} = 2 e^{\frac{i \pi }{24}} \]

OpenStudy (anonymous):

this is what my teacher wrote (16cis30)^1/4 2cis(7.5+90k) , k=0,1,2,3

OpenStudy (anonymous):

k=0: 2cis7.5 k=1: 2cis97.5 k=2: 2cis187.5 k=3: 2cis277.5

OpenStudy (anonymous):

You gave the general answer. That is great.

OpenStudy (anonymous):

i still dont get it thow i just wrote down what my teacher wrote on the board xD

OpenStudy (anonymous):

how did he get 7.5 in 2cis(7.5+90k)?

OpenStudy (anonymous):

oh wait i get it he divided 30 by 4 :)

OpenStudy (anonymous):

but why would i add 7.5 by 90

OpenStudy (anonymous):

in 2cis(7.5+90k)

OpenStudy (anonymous):

Every non zero complex number has n n_th roots. \[ (8\sqrt 3 + 8 i)^{\frac 1 4} = \left(16 e^{ i \left(\frac \pi 6 + 2 k \pi\right)}\right)^{\frac 1 4}= 2 e^{ i \left(\frac \pi {24} + k \frac \pi 2\right)}, \, k=0, 1, 2 ,3 \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

so that's why i add it by 90?